A solution of salt water was made by dissolving 4.25 g of salts for every liter of water. If the salt water solution consists of only NaCl, calculate the osmotic pressure of the salt water solution at 35°C. (Hint: NaCl is an electrolyte)
Answer #1

Ans. Osmotic pressure = Vant hoff factor× Conc of solution× gas constant× temperature

  • Vant hoff factor = 2 since there are two ions Na+ and Cl-.
  • Concentration of solution is 4.25÷40 = 0.10625 M
  • Gas constant = 0.0821 L-atm/mol-K
  • Temperature = 35+ 273= 308K

Osmotic pressure = 2× 0.10625× 0.0821×308

=5.37 atm

Please give a thumbs up??? if it helped you?